toms uncle owns a triangular piece of land. the perimeter fence that surrounds the land measures 378 yards. the shortest side is 30 yards longer than one-half of the longest side. the second longest side is 2 yards shorter than the longest side. what is the length of each side?: toms uncle owns a triangular piece of land. the perimeter fence that surrounds the land measures 378 yards. the shortest side is 30 yards longer than one-half of the longest side. the second longest side is 2 yards shorter than the longest side. what is the length of each side?
Answer by ankor@dixie-net.com(4538)

the fence measures 378 yards. the shortest side is 30 yards longer than one-half of the longest side. the second longest side is 2 yards shorter than the longest side. what is the length of each side?
:
Let x = the longest side
:
from the statements,
"the shortest side is 30 yards longer than one-half of the longest side."
we can say:
.5x + 30 = the shortest side
and
" the second longest side is 2 yards shorter than the longest side."
x-2 = 2nd longest side
;
what is the length of each side?
:
Solve for x:
x + (.5x+30) + (x-2) = 378
x + .5x + x + 30 - 2 = 378
2.5x + 28 = 378
2.5x = 378 - 28
2.5x = 350
x = 350/2.5
x = 140 yds is the longest side
then
.5(140) + 30 = 100 yd is the shortest side
and
140 - 2 = 138 yds is the 2nd side
:
:
Check: 140 + 100 + 138 = 378